# Training-free LLM Verification via Recycling Few-shot Examples

### v1

@@ Abstract
Although LLMs have achieved remarkable performance, the inherent stochasticity of their reasoning process and varying conclusions present significant challenges. Majority voting or Best-of-N with external verification models has been explored to find the most promising solution among multiple LLM outputs. However, these approaches have certain limitations, such as limited applicability or the cost of an additional training step. To address this problem, we propose a novel and effective framework that Recycles Few-shot examples to verify LLM outputs (ReFeri). Our key idea is to additionally utilize the given few-shot examples to evaluate the candidate outputs of the target query, not only using them to generate outputs as the conventional few-shot prompting setup. Specifically, ReFeri evaluates the generated outputs by combining two different scores, designed motivated from Bayes’ rule, and subsequently selects the candidate that is both confidently determined and contextually coherent through a few additional LLM inferences. Experiments with three different LLMs and across seven diverse tasks demonstrate that our framework significantly improves the accuracy of LLMs-achieving an average gain of 4.8%-through effective response selection, without additional training. [^1] Code is available at https://github.com/dongseok1220/Referi.
**footnotetext: This work was done while the authors was an intern at Yonsei ML3.
@@ 1 Introduction
Recently, large language models (LLMs) have shown remarkable performance in many real-world tasks involving complex reasoning such as math, coding, and robotics [1, 2, 3, 4]. To enhance the reasoning capacity of LLMs, various approaches have been proposed from in-context learning at test time [5, 6] to recent RL training method [7, 8]. Despite these improvements, the inherent stochastic nature of LLM still presents significant challenges, since different reasoning paths can be generated for the same input and can lead to varying conclusions [9, 10, 11]. Majority voting approaches, such as self-consistency [12, 13], have been widely adopted to reduce such randomness by aggregating multiple LLM outputs and determining a single prediction. However, this approach is only applicable when the answer can be easily extracted from the output and aggregated. Consequently, it is difficult to apply to open-ended text generation tasks such as summarization and personalized chatbot [14, 15].
@@
To address this challenge, finding the most promising one among multiple LLM outputs using a specific selection method, often called Best-of-N, has recently gained attention [16, 17]. For instance, one of the most representative approaches is to score each output using external verification models such as Outcome Reward Models (ORMs) [18, 19] or Process Reward Models (PRMs) [20, 21], and then selecting the highest-scoring output. However, to obtain these reward models, training with a large amount of task-specific labeled data is often necessary; therefore, applying this framework to specific target domain, which is far from well-explored domains such as math and coding, is challenging. Prompting LLM to select the most promising output such as LLM-as-judge is considerable to remove the reliance on the verification model [22, 23]. However, this approach is only effective when the given LLM has sufficient intrinsic knowledge for the target domain; consequently, it often requires separate training steps and datasets again to achieve sufficient performance [24, 25, 26].
@@
Motivated by this, we suggest a new perspective: utilization of few-shot examples to verify and select among multiple LLM outputs. As recent LLMs have been trained with an extensive instruction tuning step, they often exhibit better performance without few-shot examples [8, 27], and hence using these examples at test time is recently losing attention (see results in Table 1). However, we argue that using few-shot examples is still one of the easiest and most direct ways to let LLMs know how to solve the given task with human prior knowledge, even if LLMs have not encountered it before. Therefore, in this work, we provide a new framework that enables better exploitation of few-shot examples by using them not only for generating multiple outputs, but also for selecting the most promising one.
@@
Figure 1: An overview of ReFeri. For $K$ candidate responses from LLMs, ReFeri assigns each candidate a forward confidence score (how likely candidate is to be generated conditioned on few-shot examples) and a backward consistency score (how candidate affect to reconstruct the answers of few-shot examples). Then, the response with the best joint score is selected as the final answer.
@@
Contribution. In this work, we propose ReFeri, a novel and effective framework that Recycles Few-shot examples to verify LLM outputs. The core idea of ReFeri is additionally utilizing the given few-shot examples to evaluate the candidate outputs of the target query, not only using them to generate outputs as conventional few-shot in-context learning. [^2] In-context learning uses given few-shot examples as additional input context upon the target query. Specifically, ReFeri estimates the likelihood of the generated outputs by decomposing it into two different scores conditioned on few-shot examples, which are derived from Bayes’ rule. The forward confidence score measures the likelihood of candidate outputs given the few-shot examples and the test query, favoring more confident ones. On the other hand, the backward consistency score measures whether conditioning on the candidate output affects the likelihood of the few-shot examples compared to conditioning on their queries alone. By combining these scores, ReFeri selects the candidate that is both confidently determined and contextually coherent through a few additional LLM inferences. Consequently, ReFeri does not require additional model training to select the most promising output, and allows better leverage of both intrinsic knowledge of LLM and human prior within the provided few-shot examples. See Figure 1 for the illustration.
@@
We validate the effectiveness of ReFeri across three different LLMs (GPT-4o, GPT-4o-mini, and LLaMA-3.1-8B) and seven different benchmarks. When selecting one response among five candidates generated by few-shot chain-of-thought(CoT) prompting, ReFeri consistently outperforms other training-free selection across all tasks, with an average gain of 4.8% over random selection and 2.6% over prompt-based selection methods (see Figure 2). ReFeri also scales reliably with the number of candidate responses, demonstrating its practical utility in test-time scaling. To better understand the behavior of ReFeri, we conduct two complementary analyses, showing that our method is robust to both variations in few-shot example selection and the choice of model used for likelihood estimation. Furthermore, we evaluate ReFeri in more challenging and practical scenarios. Although recent reasoning LLMs [8] show degraded performance under a few-shot prompting, ReFeri consistently finds out the most promising reasoning path. Moreover, ReFeri shows effectiveness on LLM personalization [15] requires open-ended, user-specific generation.
2 Related Works
Few-shot in-context learning of LLM. Few-shot in-context learning (ICL) revealed that LLMs can generalize to unseen tasks with just a handful of input-output demonstrations [28]. To handle complex reasoning problems, chain-of-thought (CoT) prompting was proposed to append intermediate steps to the few-shot examples, leading to substantial gains in tasks such as arithmetic, commonsense reasoning, and symbolic manipulation [5, 29, 30]. To further enhance ICL, various strategies have been developed to retrieve better examples using semantic similarity or entropy-based selection [31, 32]. However, some studies have shown that few-shot ICL does not always guarantee improvements. For instance, label shuffling or format changes can often leave performance unaffected [33], and the performance gap between zero-shot and few-shot CoT is narrowing in several benchmarks as instruction tuning becomes more effective [27]. In particular, recent LLMs such as DeepSeek-R1, which are trained with reinforcement learning-based reasoning steps, sometimes even show performance degradation when few-shot CoT examples are added [8]. Nonetheless, carefully selected demonstrations are still effective [34]. For example, [35] show that few-shot examples can reduce overconfidence in multi-step reasoning, and [36] show that they help mitigate hallucinations and memory-based mistakes in complex tasks. These observations motivate us to go beyond using few-shot examples solely for generation, and recycling few-shot examples to evaluate the possible multiple LLM responses and to select the most promising one.
Selection of diverse LLM outputs. Due to the probabilistic nature of LLM decoding, LLM can provide diverse outputs for a single input, each reflecting different reasoning paths [9, 10, 11]. To handle this variability, self-consistency [12] samples $K$ independent reasoning paths and selects the majority answer to improve accuracy. However, it assumes that the model produces a single, well-formatted answer, and this assumption is often violated in open-ended tasks such as summarization or free-form dialogue [14, 15]. Alternatively, recent Best-of-N approaches aim to directly select the best output among candidates, often using external verification models. For instance, Outcome Reward Models (ORMs) grade final outputs [18, 19], while Process Reward Models (PRMs) assess intermediate reasoning steps to provide finer supervision [20, 21]. Despite their successes, these models require large-scale, task-specific annotations or domain-specific checkers, limiting their scalability to new domains or unseen tasks. To eliminate the need for external verification models, prompting-based methods such as LLM-as-Judge ask LLM to evaluate its own outputs [22, 23]. However, their effectiveness heavily depends on the model’s prior knowledge in the target domain. When this knowledge is lacking, these methods require additional fine-tuning with curated evaluation datasets for sufficient performance, which reintroduces the need for supervision [24, 25, 26]. In contrast, ReFeri is training-free and task-agnostic by recycling a few-shot examples for verification.
3 Training-free Verification of LLM Outputs via Recycling Few-shot Data
3.1 Preliminary
@@
where multiple non-identical predictions $r_{k},k=1,\dots,K$ can be sampled. Then, our goal is to find the most appropriate response $r_{k^{*}}$ among them. For example, the self-consistency method [12] simply applies majority voting to determine the single prediction. On the other hand, the best-of-K method uses the external verifier such as reward models [18, 20] to score the predictions and select the highest scored one. Formally, with the external verifier $R_{\phi}$, it can be described as below:
@@
3.2 ReFeri: Verification of LLM outputs with Bayes-inspired scores with few-shot data
@@
Problem setup. Let us assume that we have an estimation model $P$ which can measure the likelihood $P(\mathbf{y}_{k})=P(r_{k}\mid\widetilde{q})$ of the response $r_{k}$ conditioned on the given query $\widetilde{q}$. [^3] For the experiments in Section 4, we use pre-trained LLM as the estimation model. Then, our goal is to select the response $r_{k^{*}}$ which yields the highest likelihood if the estimation is accurate:
@@
We note that the likelihood has shown effectiveness to find high-quality reasoning path [10]. However, selecting based on the estimated $P(\mathbf{y}_{k})$ could be ineffective in practice, as it entirely depends on the estimation model’s intrinsic knowledge, which can be limited in unfamiliar or challenging domains. Furthermore, when there is a mismatch between $\mathcal{M}$ and $P$, the estimated likelihoods can be unreliable as minor syntactic variations in response can make large deviations. To address this, we propose to reinterpret $P(\mathbf{y_{k}})$ with few-shot examples $\mathbf{X}$, through Bayes’ rule:
@@
Backward consistency score. The backward score, $\log P(\mathbf{X}|\mathbf{y}_{k})-\log P(\mathbf{X})$, evaluates how the inclusion of test query $\widetilde{q}$ and candidate response $r_{k}$ affect to explain the few-shot examples $\mathbf{X}$. At a high level, this score serves as a form of consistency check between the response and the given few-shot examples. Under the assumption of mutual independence between few-shot examples, the backward score can also be derived similar to Eq. 6. However, to better evaluate how well the candidate response $r_{k}$ explains the few-shot examples $\mathbf{X}$, we refine the backward term using a leave-one-out strategy [37, 38] through prompt replacement; namely, we construct new demonstration $\widetilde{\mathbf{X}}_{i}$ by replacing $i$-th example $\mathbf{x}_{i}=(q_{i},a_{i})$ with a pair of test query and candidate response $(\widetilde{q},r_{k})$:
@@
$S_{\tt Back}(\mathbf{y_{k}}):=\log P(\mathbf{X}\mid\mathbf{y}_{k})-\log P(\mathbf{X})=\sum_{i=1}^{N}\big{(}\log P(a_{i}\mid q_{i},\widetilde{\mathbf{X}}_{i})-\log P(a_{i}\mid q_{i})\big{)}.$ (8)
@@
Final score. By combining forward and backward scores following Eq. 5, we design our main selection score $S_{\tt Comb}$ to find the most promising output $r_{k^{\star}}$ as below:
${k^{\star}}=\arg\max_{k=1,\dots,K}S_{\tt Comb}(\mathbf{y_{k}}),\quad S_{\tt Comb}(\mathbf{y_{k}}):=S_{\tt Forw}(\mathbf{y_{k}})-S_{\tt Back}(\mathbf{y_{k}}).$ (9)
However, the direct estimation of $P(\mathbf{y}_{k})$ can be effective, particularly when both the estimation model $P$ and the generation LLM $\mathcal{M}$ perform reliably—as is often the case in well-covered, generic domains. Therefore, to utilize this information upon the proposed combined score $S_{\tt Comb}$, we define the final score $S_{\tt Fin}$ by jointly considering both components as follows:
$S_{\tt Fin}(\mathbf{y_{k}}):=(1-\beta)\cdot S_{\tt Dire}(\mathbf{y_{k}})+\beta\cdot S_{\tt Comb}(\mathbf{y_{k}}),$ (10)
where $S_{\tt Dire}=P(\mathbf{y}_{k})$ and $\beta$ is a hyperparameter to control the effect from two sources. This is helpful to prevent over-reliance on noisy few-shot examples of $S_{\tt Comb}$ and improves the robustness of selection. When $\beta=1$, the score relies fully on the score derived forward-backward decomposition.
@@ 4 Experiments
    What is the effect of each component in ReFeri? (Tables 2, 3, and Figure 4)
@@
    Can ReFeri be generalized to recent reasoning LLM and broader tasks? (Tables 4, 5)
@@ 4.1 Setups
Datasets. We evaluate our method on seven benchmarks encompassing diverse reasoning paradigms, including symbolic-numeric, expertise-based, and multi-hop textual reasoning tasks. (1) MATH500 [20]; a 500-problem subset of the MATH benchmark [39], focused on symbolic manipulation and multi-step mathematical reasoning. (2) MMLU-pro [40]; 4200 examples, including 300 randomly sampled questions per domain (e.g., physics, law, computer science) extends the original MMLU benchmark [41] by adding reasoning-focused questions and expanding the choice set from four to ten options. (3) HotpotQA [42]; 500 samples from [43] which is a multi-hop question-answering benchmark requiring reasoning across multiple Wikipedia paragraphs with annotated supporting facts. (4) DROP [44]; 500 randomly sampled questions from this reading comprehension benchmark, which queries demand discrete numerical reasoning (e.g., addition, counting, sorting) over paragraphs. (5) GPQA-diamond [45] (GPQA); 198 graduate-level questions assessing complex reasoning in biology, physics, and chemistry. (6,7) MuSR [46]; 256 examples in Object Placement (MuSR-op) and 250 examples in Team Allocation (MuSR-ta) tasks assessing spatial and relational reasoning.
@@
Notably, prior work [27] has shown that few-shot Chain-of-Thought (CoT) prompting yields significant gains over zero-shot CoT in MuSR, highlighting the role of in-context examples in complex reasoning. As few-shot examples are necessary for some baselines and ReFeri, we collect them following the previous works. MATH500: 5 examples from [47] (GPTs), 4 examples from [48] (LLaMA).444(1) Using the same prompt as GPT results in significantly lower accuracy, and (2) LLaMA-based models provide their own optimized prompt templates (see meta-llama/Llama-3.2-3B-Instruct-evals). MMLU-Pro: 5 examples from [40]. HotpotQA: 6 examples from [49]. DROP: 3 examples following [50]. GPQA-Diamond: 5 examples from [45]. MuSR: 3 examples from [27]. Complete prompt templates are available in Appendix A.1.
Baselines.
@@
We compare ReFeri against five prompt-based methods that require no additional training, with some reflecting different uses of few-shot examples: (1) Zero-shot CoT appends a trigger phrase ("Let’s think step by step.") to each query without providing exemplars, relying on LLM’s intrinsic reasoning capabilities. (2) Few-shot CoT prepends a fixed set of few examples, enabling LLM to generalize from few in-context demonstrations. (3) USC asks LLM to select the best answer from multiple CoT outputs, by following [22]. We report the results without using few-shot examples as it degrades performance (see Appendix B.3). (4) CoT-WP [10] scores each candidate response using token-level probabilities from LLM conditioned on the same few-shot examples. Specifically, the score is a confidence gap between top-1 and top-2 tokens at answer positions. To estimate confidence, we use LLaMA-3.1-8B as same as ReFeri. (5) LEAP [51] improves few-shot prompting by intentionally inducing mistakes on few examples. Then extracting generalizable task-specific principles through self-reflection without human annotations, and prompting the model to apply these principles to unseen questions. Specific prompts for each baseline are in Appendix A.2.
@@
Implementation details. For the experiments, we use (1) gpt-4o-2024-08-06 (GPT-4o) [52], (2) gpt-4o-mini-2024-07-18 (GPT-4o-mini) [53], and (3) LLaMA-3.1-8B-Instruct (LLaMA-3.1-8B) [2] as target LLMs, i.e., response generation models. We generate $K=5$ responses per each query using temperature of $1.0$ to encourage diverse candidates. For Zero-shot CoT, Few-shot CoT and LEAP, we report the average accuracy for all five responses without applying any selection mechanism, which can be viewed as randomly selecting the response. For USC, COT-WP, and ReFeri, we use the same candidates generated from Few-shot CoT. In USC, the generation and judge models are same and we use temperature of 00. For the estimation model $P$, we employ LLaMA-3.1-8B-Instruct, except the experiments in Figure 4. For the hyper-parameter of ReFeri, we consider $\beta\in[0.5,0.75,1]$. More details are in Appendix A.3.

### v2

@@ Abstract
Although LLMs have achieved remarkable performance, the inherent stochasticity of their reasoning process and varying conclusions present significant challenges. Majority voting or Best-of-N with external verification models has been explored to find the most promising solution among multiple LLM outputs. However, these approaches have certain limitations, such as limited applicability or the cost of an additional training step. To address this problem, we propose a novel and effective framework that Recycles Few-shot examples to verify LLM outputs (ReFeri). Our key idea is to additionally utilize the given few-shot examples to evaluate the candidate outputs of the target query, not only using them to generate outputs as the conventional few-shot prompting setup. Specifically, ReFeri evaluates the generated outputs by combining two different scores, designed motivated by Bayes’ rule, and subsequently selects the candidate that is both confidently determined and contextually coherent through a few additional LLM inferences. Experiments with three different LLMs and across seven diverse tasks demonstrate that our framework significantly improves the accuracy of LLM–achieving an average gain of 4.5%–through effective response selection, without additional training.
@@ 1 Introduction
Recently, large language models (LLMs) have shown remarkable performance in many real-world tasks involving complex reasoning, such as math, coding, and robotics (Anthropic, 2024; Dubey et al., 2024; OpenAI, 2024c; Team et al., 2023). To enhance the reasoning capacity of LLMs, various approaches have been proposed, ranging from in-context learning at test time (Wei et al., 2022; Kojima et al., 2022) to recent RL training method (Qu et al., 2024; Guo et al., 2025). Despite these improvements, the inherent stochastic nature of LLM still presents significant challenges, since different reasoning paths can be generated for the same input and can lead to varying conclusions (Kadavath et al., 2022; Wang & Zhou, 2024; Qiu & Miikkulainen, 2024). Majority voting approaches, such as self-consistency (Wang et al., 2023b; Aggarwal et al., 2023), have been widely adopted to reduce such randomness by aggregating multiple LLM outputs and determining a single prediction. However, this approach is only applicable when the answer can be easily extracted from the output and aggregated. Consequently, it is difficult to apply to open-ended text generation tasks such as summarization and personalized chatbot (Stiennon et al., 2020; Salemi et al., 2024).
@@
To address this challenge, finding the most promising one among multiple LLM outputs using a specific selection method, often called Best-of-N, has recently gained attention (Snell et al., 2024; Gui et al., 2024). For instance, one of the most representative approaches is to score each output using external verification models such as Outcome Reward Models (ORMs) (Cobbe et al., 2021; Uesato et al., 2022) or Process Reward Models (PRMs) (Lightman et al., 2024; Wang et al., 2024a), and then selecting the highest-scoring output. However, to obtain these reward models, training with a large amount of task-specific labeled data is often necessary; therefore, applying this framework to specific target domain, which is far from well-explored domains such as math and coding, is challenging. Prompting LLM to select the most promising output–such as LLM-as-judge–is considerable to remove the reliance on the verification model (Chen et al., 2023; Zheng et al., 2023). However, this approach is only effective when the given LLM has sufficient intrinsic knowledge for the target domain; consequently, it often requires separate training steps and datasets again to achieve sufficient performance (Yuan et al., 2024; Mahan et al., 2024; Zhang et al., 2025).
@@
Motivated by this, we suggest a new perspective: utilization of few-shot examples to verify and select among multiple LLM outputs. As recent LLMs have been trained with an extensive instruction tuning step, they often exhibit better performance without few-shot examples (Guo et al., 2025; Sprague et al., 2025), and hence using these examples at test time is recently losing attention (see results in Table 1). However, we argue that using few-shot examples is still one of the easiest and most direct ways to let LLMs know how to solve the given task with human prior knowledge, even if LLMs have not encountered it before. Therefore, in this work, we provide a new framework that enables better exploitation of few-shot examples by using them not only for generating multiple outputs, but also for selecting the most promising one.
@@
Figure 1: An overview of ReFeri. For $K$ candidate responses from LLMs, ReFeri assigns each candidate a forward confidence score (how likely candidate is to be generated conditioned on few‑shot examples) and a backward consistency score (how well candidate explains the answers of few‑shot examples). Then, the response with the best joint score is selected as the final answer.
@@
Contribution. In this work, we propose ReFeri, a novel and effective framework that Recycles Few-shot examples to verify LLM outputs. The core idea of ReFeri is additionally utilizing the given few-shot examples to evaluate the candidate outputs of the target query, not only using them to generate outputs as conventional few-shot in-context learning. [^1] In-context learning uses given few-shot examples as additional input context upon the target query. Specifically, ReFeri estimates the likelihood of the generated outputs by decomposing it into two different scores (forward confidence score and backward consistency score) conditioned on few-shot examples, which are derived from Bayes’ rule. The forward confidence score measures the likelihood of candidate outputs given the few-shot examples and the test query, favoring more confident ones. On the other hand, the backward consistency score measures whether conditioning on the candidate output well explains the likelihood of the few-shot examples compared to conditioning on their queries alone. By combining these scores, ReFeri selects the candidate that is both confidently determined and contextually coherent through a few additional LLM inferences. Consequently, ReFeri does not require additional model training to select the most promising output, and allows better leverage of both intrinsic knowledge of LLM and human prior within the provided few-shot examples. See Figure 1 for the illustration.
@@
We validate the effectiveness of ReFeri across three different LLMs (GPT‑4o, GPT‑4o‑mini, and LLaMA‑3.1‑8B) and seven different benchmarks. When selecting one response among five candidates generated by few-shot chain-of-thought (CoT) prompting, ReFeri consistently outperforms other training-free selection across all tasks, with an average gain of 4.5% over random selection and 2.4% over prompt-based selection methods (see Figure 2). ReFeri also scales reliably with the number of candidate responses, demonstrating its practical utility in test-time scaling. To better understand the behavior of ReFeri, we conduct more complementary analyses, showing that our method is robust to variations in few-shot example selection, prompt template choices, and the choice of model used for likelihood estimation; ReFeri yields consistent improvements without reliance on specific prompt templates or few-shot examples. Moreover, when combined with smaller likelihood estimators, ReFeri significantly reduces the cost per query compared to larger baseline models while outperforming them.
2 Training-free LLM Verification via Recycling Few-shot Data
2.1 Problem formulation
@@
where multiple non-identical predictions $r_{k},k=1,\dots,K$ can be sampled. Then, our goal is to find the most appropriate response $r_{k^{*}}$ among them. For example, the self-consistency method (Wang et al., 2023b) simply applies majority voting to determine the single prediction. On the other hand, the best-of-K method uses the external verifier such as reward models (Cobbe et al., 2021; Lightman et al., 2024) to score the predictions and select the highest scored one. Formally, with the external verifier $R_{\phi}$, it can be described as below:
@@
2.2 ReFeri: Verifying LLM outputs with Bayes’ rule with few-shot data
@@
[…]
@@
We note that the likelihood has shown effectiveness to find high-quality reasoning path (Wang & Zhou, 2024). However, selecting based on the estimated $P(\mathbf{y}_{k})$ could be ineffective in practice, as it entirely depends on the estimation model’s intrinsic knowledge, which can be limited in unfamiliar or challenging domains. Furthermore, when there is a mismatch between $\mathcal{M}$ and $P$, the estimated likelihoods can be unreliable as minor syntactic variations in response can make large deviations. To address this, we propose to reinterpret $P(\mathbf{y_{k}})$ with few-shot examples $\mathbf{X}$, through Bayes’ rule:
@@
Backward consistency score. The backward score, $\log P(\mathbf{X}|\mathbf{y}_{k})-\log P(\mathbf{X})$, evaluates how well the test query $\widetilde{q}$ and candidate response $r_{k}$ explains the few-shot examples $\mathbf{X}$. At a high level, this score serves as a form of consistency check between the response and the given few-shot examples. Under the assumption of mutual independence between few-shot examples, the backward score can also be derived similar to Eq. 6. To better utilize given few-shot examples, we refine the backward term using a leave-one-out strategy (Perez et al., 2021; Izacard et al., 2023) through prompt replacement; namely, we construct new demonstration $\widetilde{\mathbf{X}}_{i}$ by replacing $i$-th example $\mathbf{x}_{i}=(q_{i},a_{i})$ with a pair of test query and candidate response $(\widetilde{q},r_{k})$:
@@
$S_{\tt Back}(\mathbf{y_{k}}):=\log P(\mathbf{X}\mid\mathbf{y}_{k})-\log P(\mathbf{X})=\frac{1}{N}\sum_{i=1}^{N}\big(\log P(a_{i}\mid q_{i},\widetilde{\mathbf{X}}_{i})-\log P(a_{i}\mid q_{i})\big).$ (8)
@@
While the backward consistency score provides a reliable consistency signal, one may concern its computational cost as it grows linearly with the number of few-shot examples. To alleviate this, we propose a lightweight approximation; instead of iterating over all $N$ few-shot examples, we select only the single most relevant example to the test query $\widetilde{q}$. Specifically, we employ a pre-trained embedding model E{E} to encode both $\widetilde{q}$ and each $q_{i}$, and identify the most relevant example $\mathbf{x}_{i^{\dagger}}$:
$i^{\dagger}=\arg\max_{i=1,\dots,N}\cos\big(E(\widetilde{q}),E(q_{i})\big).$ (9)
Then, we define the approximated backward score $\widetilde{S}_{\tt Back}$:
$\widetilde{S}_{\tt Back}(\mathbf{y_{k}}):=\log P(a_{i^{\dagger}}\mid q_{i^{\dagger}},\widetilde{\mathbf{X}}_{i^{\dagger}})-\log P(a_{i^{\dagger}}\mid q_{i^{\dagger}})$ (10)
Final score. By combining forward and backward scores following Eq. 5, we design our main selection score $S_{\tt Fin}$ to find the most promising output $r_{k^{\star}}$ as below:
${k^{\star}}=\arg\max_{k=1,\dots,K}S_{\tt Fin}(\mathbf{y_{k}}),\quad S_{\tt Fin}(\mathbf{y_{k}}):=S_{\tt Forw}(\mathbf{y_{k}})-\widetilde{S}_{\tt Back}(\mathbf{y_{k}}).$ (11)
@@ 3 Experiments
    What is the effect of each component, and how robust is ReFeri? (Tables 2, 3, 4)
@@
    How does the estimation model affect cost and performance of ReFeri? (Figure 4, Table 15)
@@ 3.1 Setups
Datasets. We evaluate our method on seven benchmarks encompassing diverse reasoning paradigms, including symbolic-numeric, expertise-based, and multi-hop textual reasoning tasks. (1) MATH500 (Lightman et al., 2024); a 500-problem subset of MATH (Hendrycks et al., 2021b), focused on symbolic manipulation and multi-step mathematical reasoning. (2) MMLU-pro (Wang et al., 2024b); 4200 examples, including 300 randomly sampled questions per domain (e.g., physics, law, computer science) extends the original MMLU benchmark (Hendrycks et al., 2021a) by adding reasoning-focused questions and expanding the choice set from four to ten. (3) HotpotQA (Yang et al., 2018); 500 samples from (Kim et al., 2024) a multi‑hop question‑answering benchmark requiring reasoning across multiple Wikipedia paragraphs with annotated supporting facts. (4) DROP (Dua et al., 2019); 500 randomly sampled questions from this reading comprehension benchmark, demanding discrete numerical reasoning (e.g., addition, counting, sorting) over paragraphs. (5) GPQA-diamond (Rein et al., 2024) (GPQA); 198 graduate-level questions assessing complex reasoning in biology, physics, and chemistry. (6,7) MuSR (Sprague et al., 2024); 256 examples in Object Placement (MuSR-op) and 250 examples in Team Allocation (MuSR-ta) tasks assessing spatial and relational reasoning.
@@
Notably, prior work (Sprague et al., 2025) has shown that few-shot Chain-of-Thought (CoT) prompting yields significant gains over zero-shot CoT in MuSR, highlighting the role of in-context examples in complex reasoning. As few-shot examples are necessary for some baselines and ReFeri, we collect them following the previous works. MATH500: 5 examples from (Yang et al., 2024) (GPTs), 4 examples from (Lewkowycz et al., 2022) (LLaMA).333(1) Using the same prompt as GPT results in significantly lower accuracy, and (2) LLaMA-based models provide their own optimized prompt templates (see meta‑llama/Llama‑3.2‑3B‑Instruct‑evals). MMLU-Pro: 5 examples from (Wang et al., 2024b). HotpotQA: 6 examples from (Yao et al., 2023). DROP: 3 examples following (Zhou et al., 2022). GPQA-Diamond: 5 examples from (Rein et al., 2024). MuSR: 3 examples from (Sprague et al., 2025). Complete prompt templates are available in Appendix A.1.
@@
Baselines. We compare ReFeri against five widely-used prompt-based methods that require no additional training, with some reflecting different uses of few-shot examples: (1) Zero-shot CoT appends a trigger phrase ("Let’s think step by step.") to each query without providing exemplars, instead relying on LLM’s intrinsic reasoning capabilities. (2) Few-shot CoT prepends a fixed set of few examples, enabling LLM to generalize from few in-context demonstrations. (3) LEAP (Zhang et al., 2024) improves few-shot prompting by intentionally inducing mistakes on few examples. Then extracting generalizable task-specific principles through self-reflection without human annotations, and prompting the model to apply these principles to unseen questions. Specific prompts for each baseline are in Appendix A.2. (4) USC asks LLM to select the best answer from multiple CoT outputs, by following (Chen et al., 2023). (5) CoT-WP (Wang & Zhou, 2024) scores each candidate response using token-level probabilities from LLM conditioned on the same few-shot examples. Specifically, the score is a confidence gap between top-1 and top-2 tokens at answer positions.
@@
Implementation details. For the experiments, we use (1) gpt-4o-2024-08-06 (GPT-4o) (OpenAI, 2024a), (2) gpt-4o-mini-2024-07-18 (GPT-4o-mini) (OpenAI, 2024b), and (3) LLaMA-3.1-8B-Instruct (LLaMA-3.1-8B) (Dubey et al., 2024) as target LLMs, i.e., response generation models. We generate $K=5$ responses per query using temperature of $1.0$ to encourage diverse candidates. For Zero-shot CoT, Few-shot CoT and LEAP, we report the average accuracy across five responses without applying any selection mechanism, which can be viewed as randomly selecting the response. For USC, CoT-WP, and ReFeri, we use the same candidates generated from Few-shot CoT and employ LLaMA-3.1-8B-Instruct as the estimation (or LLM-judge) model, except in the experiments Figure 4 and Table 15. In USC (i.e., LLM-as-Judge setting), the decoding temperature is fixed at 0 for determinism. For computing similarity in backward consistency score (Eq. 9), we employ the lightweight embedding model all-mpnet-base-v2 with 110M parameters. More details are in Appendix A.3.